OpenStudy (otonogold):

Multiple Choice which of the following is equal to sqrt 18 x^9y^4? A. 3x^3y^2 B. 3 x^4y^2 sqrt 2x C. 9 x^3y^2 D. 3 x^3y^2 sqrt 2

OpenStudy (otonogold):

@jim_thompson5910

OpenStudy (otonogold):

I think its A or D

OpenStudy (otonogold):

@agent0smith

jimthompson5910 (jim_thompson5910):

I'll get you started \[\Large \sqrt{18x^9y^4}\] \[\Large \sqrt{9*2*x^9*y^4}\] \[\Large \sqrt{9*2*x^{8+1}*y^4}\] \[\Large \sqrt{9*2*x^8*x^1*y^4}\] \[\Large \sqrt{9*2*x^{4*2}*x^1*y^{2*2}}\] \[\Large \sqrt{9*x^{4*2}*y^{2*2}*2*x}\] \[\Large \sqrt{3^2*(x^4)^2*(y^2)^2*2*x}\] \[\Large \sqrt{(3x^4y^2)^2*2x}\] And I'll let you finish up

OpenStudy (otonogold):

why would x^4 and not x^3

OpenStudy (agent0smith):

Since you have an (x^4)^2 and an x, which all together gives x^9

OpenStudy (otonogold):

being 3 x^4y^2 sqrt 2x. Thank you, I figure it out

OpenStudy (agent0smith):

^yep